Klumper Corporation is a diversified manufacturer of industrial goods. The company’s activity-based costing system contains the following six activity cost pools and activity rates:

Activity Cost Pool Activity Rates
Supporting direct labor $ 11 per direct labor-hour
Machine processing $ 5 per machine-hour
Machine setups $ 45 per setup
Production orders $ 160 per order
Shipments $ 125 per shipment
Product sustaining $ 800 per product

Activity data have been supplied for the following two products:

Total Expected Activity

K425 M67
Number of units produced per year 200 2,000
Direct labor-hours 975 30
Machine-hours 3,000 20
Machine setups 5 3
Production orders 5 3
Shipments 10 3
Product sustaining 3 3

Required:

How much total overhead cost would be assigned to K425 and M67 using the activity-based costing system?

Activity Cost Pool K425 M67
Supporting direct labor
Machine processing
Machine setups
Production orders
Shipments
Product sustaining
Total overhead cost
Answer #1
K425 M67
Supporting direct labor $11*975 = $10,725 $11*30 = $330
Machine processing $5*3,000 = $15,000 $5*20 = $100
Machine setups $49*5 = $245 $49*3 = $147
Production orders $160*5 = $800 160*3 = $480
Shipments $125*10 = $1,250 $125*3 = $375
Product sustaining $800*3 = $2,400 $800*3 = $2,400
Total overhead cost $                        30,420 $                    3,832

You can reach me over comment box if you have any doubts. Please rate this answer